OpenStudy (anonymous):

Solve for x by completing the square. 9x^2+9x−4=0

OpenStudy (anonymous):

You just need to factor the given equation. In this case: 9x^2 - 9x - 4 = 0 (3x - 4)(3x + 1) = 0 So the possible answers are -1/3 and 4/3, since if either of those is the value of x, the value of the whole expression on the left is 0, making the statement true.